### v1

# OnlyFlow: Optical Flow based Motion Conditioning for Video Diffusion Models

Abstract
color=yellow!60,inline]H:regle du pouce, ajouter texte entre section et sous sections Dire dans les RW, si methode comparable, expliquer puis critiquer, et pourquoi notre modele a lieu d’être)
We consider the problem of text-to-video generation tasks with precise control for various applications such as camera movement control and video-to-video editing. Most methods tacking this problem rely on providing user-defined controls, such as binary masks or camera movement embeddings. In our approach we propose OnlyFlow, an approach leveraging the optical flow firstly extracted from an input video to condition the motion of generated videos. Using a text prompt and an input video, OnlyFlow allows the user to generate videos that respect the motion of the input video as well as the text prompt. This is implemented through an optical flow estimation model applied on the input video, which is then fed to a trainable optical flow encoder. The output feature maps are then injected into the text-to-video backbone model. We perform quantitative, qualitative and user preference studies to show that OnlyFlow positively compares to state-of-the-art methods on a wide range of tasks, even though OnlyFlow was not specifically trained for such tasks. OnlyFlow thus constitutes a versatile, lightweight yet efficient method for controlling motion in text-to-video generation.
Models and code will be made available on GitHub and HuggingFace.
Figure 1: OnlyFlow controls the generation of video with text and the motion of a video input, using an estimation of its optical flow. We highly encourage readers to check our supplementary content for video results, which cannot be well represented by still images.
color=teal!60,inline]M: alpha 0.5 * 2 flow superposition
CVPR 2023 transfer optical flow: https://diffusion-motion-transfer.github.io/MotionEditing_arXiv.pdf
fig 2 : changer RAFT
loss de training dans related work et citer dans partie 4
alpha en parametre dans la fig 2
3 Framework / architecture la plus generale possible 3.1 pas citer directement RAFT Plus à rework dans partie training 4 Training 5 Experimental results
metric intro dans fin de partie 4
1 Introduction
There has been tremendous progress in generative AI thanks to the rise of diffusion models for text-to-image (T2I) with StableDiffusion [10]. More recently T2I have been extended to text-to-video (T2V) models (such as AnimateDiff [14]). This has started revolutionizing the creation industry (entertainment professional, advertising, art) by offering increased productivity and creativity. Text-based models do not offer precise user control, as the text modality often limit the user expression. Image-to-video models [20, 15, 7] help to condition the start of the video but do not provide long range temporal conditioning. Many recent work aims to provide more precise control with the use of external conditioning data for T2I [43, 25, 41] and T2V models [23, 9, 37, 42, 13]. Users can edit parts of the generated content, with the use of masks, have style customization from existing images or even have control between the conditioning of content, layout or style.
Figure 2: Overview of OnlyFlow. Inputs are i) a tokenized and encoded text prompt, ii) noisy latents for the diffusion model and iii) the optical flow of an input video. The latter is fed through a trainable optical flow encoder which outputs features maps that are injected in the diffusion U-Net. We experiment with several injection strategies, for illustration purposes we only show the injection in temporal attention layers of the U-Net. The U-Net is kept frozen during training. The output generated video matches the prompt and the auxiliary video’s motion.
The creation of videos presents a specific challenge linked to the complexity of motion synthesis. By integrating additional motion control, one can reduce the ambiguities inherent in video synthesis. This allows for better model movements and improved ability to manipulate generated content for personalized creations.
In this paper, we introduce OnlyFlow, a simple solution to control video generation from motion. This consists in passing an auxiliary video as input to copy its motion. A simple way to describe the motion in a video is the optical flow. This description of the temporal evolution of pixels is widely used in different tasks in computer vision. OnlyFlow extracts optical flow from a given video and employs it to guide the denoising diffusion process of a video generation model based on AnimateDiff using a small external trained optical flow encoder.
To summarize, (1) we propose OnlyFlow, a new motion-guided video generation model that enables motion conditioning from given reference videos; (2) we design a motion control strategy based on optical flow representation to perform motion guidance, allowing for motion transfer across videos; (3) we present metrics on its text alignment, motion fidelity, and temporal consistency. (4) we empirically validate the effectiveness and versatility of OnlyFlow in various video generation tasks, in which several experiments demonstrate its use case.
color=teal!60,inline]M: tres tres general au debut comme l’avancée des modèles de text to image et text to video + applications et sujets d’application
questionnement de controles très important et les modeles out of the box ne permettent pas ca ainsi beaucoup de modeles ont ete apportés … citer pour donner du contrôle (ce qui s’applique aux 2 type images et videos) enumerer les types de controle - edition des parties depuis un masque - prolongement de video - influencer avec un style via image (IP-adapter) - ou composition ou layout contenu
On s’interesse aux modèles de video dont pleins de méthodes on été développées pour conditionner par d’autres inputs generalement (revenir aux autres approches), comment ? Structure commune avec un modele annexe (copie pour cn, encodeur pour t2i, etc..)
user-defined conditioning (controlnet + vectorial representation of camera movement)
mais …. pbs user defined .. dev les pbs pas facilement tedious task et ne permet pas des choses simples (ex conditionné par le mouvement, ce qui nous interresse ici)
ainsi l’objet qui nous permet d’isoler le mouvement d’une vidéo est le flow (phrase bateau sur l’utilisation du flow, maintes utilisations)
on a pris inspiration des techniques qui prenenent un modele annexe et du flow pour introduire OnlyFlow.
We introduce…(à copier) ce qu’est le modèle comment il est train ses applications (text-to-video, video-to-video style transfer, personalizing..)
les contributiosns en bulletpoint - Oui, utilisation du flow pour conditiioner animatediff - ses applications Montrer framework unique pour différentes types de tâches de customization: - ok style transfer - ok pour type contrôlnet - controle de camera - Modèle en open source, qui s’integrère facilement des plugins existant de par sa nature
color=yellow!60,inline]H:Truc a garder que j’ai tej de l’intro :
A common way to inject conditioning is to make use of an external trainable model tailored for a specific task. This model can be of the same size as the main T2I/T2V model, like in the case of the ControlNet [43], or much smaller in the case of IP-Adapter [41] and T2I-Adapter [25]
While existing models like VideoPoet [24] utilize optical flow for motion control, their weights are not publicly accessible.
2 Related work
2.1 Text-to-video diffusion models
Recent advances in text-to-image (T2I) generation have been propelled by sophisticated text encoders, enabling high-quality image synthesis from textual descriptions. Building on these successes, there’s growing interest in text-to-video (T2V) generation, aiming to extend these capabilities to generate coherent video content.
To leverage powerful pre-trained image generators like Stable Diffusion [30], some approaches introduced temporal layers [14] into existing 2D architectures and fine-tuned these models on large video datasets.
VideoLDM [4, 18] introduced a motion module that uses 3D convolutions and temporal attention to capture frame-to-frame correlations effectively. AnimateDiff [14] enhanced a pre-trained T2I diffusion model by fine-tuning specialized temporal attention layers on video data, adding motion capabilities without altering the original image generation process. This approach allows for high-quality animation creation while retaining the strengths of the base T2I model. Moreover, animatediff is compatible with all existing T2I tools such as IP-Adapter [41] and Controlnet [43].
Other methods like Align-Your-Latents [5] convert T2I models into video generators by aligning noise maps across frames. Lumiere [3] improves temporal coherence by generating videos directly at the desired frame rate, bypassing temporal super-resolution modules. Techniques such as W.A.L.T [16]. and Sora [6] operate in compressed spatiotemporal latent spaces to handle video data efficiently. Addressing data scarcity, Videocrafter2 [8] learns motion patterns from low-quality videos while capturing appearance details from high-quality images. Nevertheless, these methods lack a mature ecosystem of control and edition tools, which in practice is the reason for their slow adoption.
2.2 Controllable video generation
Recent developments in controllable image generation have sparked interest in bringing similar control mechanisms to video generation. Researchers have introduced various control signals to guide the creation of videos, such as specifying the initial frame [15, 37], defining motion trajectories [37, 31, 42], focusing on particular motion regions, and controlling specific moving objects [38]. Some methods also use images as control inputs to enhance video quality and improve temporal relationship modeling [37, 24].
Another approach involves reference-based video generation, where motion from an existing real video directs the creation of new content. Techniques that integrate frame-wise depth or edge maps aim to regulate motion in a manner similar to control mechanisms used in text-to-image models [23, 43, 9, 22]. However, these methods can unintentionally introduce static features unrelated to the intended motion, leading to inconsistencies between the generated video and the desired content.
To overcome these challenges, motion-specific fine-tuning frameworks have been developed to extract distinct motion patterns from single videos or collections with similar motion [14, 39, 13]. While these methods show promise in capturing nuanced motion, they often involve complex training processes [39] or costly denoising (DDIM) process inversion [32, 23, 13] and can risk degrading the model’s performance.
Controlling camera movement during video generation is another area of focus. Some techniques involve fine-tuning models for different shot types using methods like MotionLoRA finetuning in AnimateDiff [14]. Others introduce camera representation embeddings to manipulate camera poses [17, 38], but these may be limited by the number of camera parameters they can handle or may require fine-tuning parts of the video diffusion model, which can affect the model’s generalization ability.
Overall, while significant progress has been made in controllable video generation, providing precise control over motion of intrinsic object and camera dynamics without compromising model performance remains an ongoing challenge.
color=yellow!60,inline]H: SD + LDM / LVM
AnimateDiff
T2I-Adapter
User defined text-to-video conditioning ControlNet (c’est bien mais trop lourd) + CameraCtrl + MotionCrl
Papiers recents
3 OnlyFlow framework
3.1 Diffusion model for video generation
Text-to-video generation involves creating coherent and high-quality videos conditioned on textual prompts. Due to the high dimensionality and temporal complexity of video data, recent approaches project videos into a lower-dimensional latent space using an encoder $\mathcal{E}$ [30]. Given a batch $B$ of video $\mathbf{V}\in\mathbb{R}^{F\times H\times W\times 3}$ with $F$ frames, the latent representation is obtained as $\mathbf{z}_{0}=\mathcal{E}\left(\mathbf{V}\right)\in\mathbb{R}^{F\times\frac{H}{l}\times\frac{W}{l}\times C}$, spatialy reduce by a factor $l$. A decoder $\mathcal{D}$ reconstructs the video from the latent space, yielding $\hat{\mathbf{V}}=\mathcal{D}(\mathbf{z}_{0})$.
In this latent space, denoising diffusion probabilistic models (DDPMs) [19] are employed to approximate the latent distribution of video data. The forward diffusion process gradually adds Gaussian noise to the latent variables over $T$ timesteps, producing noisy latents $\mathbf{z}_{t}$ using the formulation:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (1)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{t}$ controls the noise schedule. The reverse diffusion process aims to recover $\mathbf{z}_{0}$ by learning a neural network $\epsilon_{\theta}$ that predicts the added noise at each timestep $t$.
To effectively capture spatial and temporal dynamics, $\epsilon_{\theta}$ is typically implemented as a 3D U-Net architecture augmented with temporal convolutions and attention mechanisms. Each feature map represents a 5D tensor of $B\times F\times H\times W\times C$. This 3D U-net is factorized over space and time, meaning that each layer only operates on the space or time dimension, leveraging spatial convolutions for frame-wise feature extraction and temporal convolutions to model motion across frames:
  • spatial: each old 2D convolution layer as in the 2D U-net is extended to be space-only 3D convolution over $H\times W\times C$. Each spatial attention block remains as attention over space
  • time: a temporal attention block is added after each spatial attention block. It performs attention over $F$. The temporal attention block is important to capture good temporal coherence.
3.2 Optical flow conditioning in T2V Architecture
In this paper, we propose to modify this traditional T2V scheme by adding motion constraint. Fig. 2 illustrates an overview of our model architecture. More precisely, for enhanced controllability in video generation, we propose to condition the video generation on an optical flow signal, extracted from an auxiliary video provided to the system. We use a standard optical flow extraction $\mathcal{T}$ technique to obtain the optical flow $\mathbf{F}=\mathcal{T}(\mathbf{V})\in\mathbb{R}^{F\times H\times W\times 2}$ of the auxiliary input video, describing pixel displacements between frames, across $F$ frames.
The Architecture of OnlyFlow extends the denoising U-Net structure of a latent diffusion model by incorporating an optical flow encoder that processes optical flow input and extracts motion features at various scales. We then proceed to inject this signal into the attention layers of the U-Net. Let’s describe these two steps:
  • Motion feature extraction: the optical flow encoder $\Phi$ extracts motion features at different scales to incorporate conditioning signals in the denoising U-Net. It initially unshuffles the optical flow to increase the number of dimensions before applying a convolutional layer. Subsequently, for each block of the U-Net, it outputs optical flow representation embeddings using a series of ResNet blocks and temporal attention mechanisms. In the temporal attention layer, incoming tensors of shape $B\times F\times H\times W\times C$ are reshaped to $(B\times H\times W)\times F\times C$ before performing attention operations across the temporal axis. This is architecture similar to that used in T2I-adapter [25] and CameraCtrl [17].
  • Motion feature injection: with the multi-scale optical flow features extracted, we inject them into the U-Net of the T2V model. The whole process is summarized in Fig. 3: we combine the latent features $h_{k}$ of the k-th block of the U-Net with the control signal from flow features $C_{k}$ through element-wise addition. The combined feature is then processed by a linear layer and scaled using a parameter $\gamma$, and its output is fed directly into the first temporal attention layer of each motion module:
    ${h_{out}}_{k}=h_{k}+\gamma\times{W_{\theta_{1}}}_{k}\left(C_{k}+h_{k}\right).$ (2)
The parameter $\gamma$ effectively controls the influence of the optical flow from the input auxiliary video onto the generated video. It is the main parameter to improve or decrease the auxiliary motion injection over the generated output. We call $\gamma$ the optical flow conditioning strength.
Figure 3: Injection strategy of the encoded optical flow conditioning $C_{k}$ from the optical flow encoder into the temporal attention layers of the $k$-th block of the U-Net.
4 Implementation and training details
We describe the experimental protocol used to implement and train OnlyFlow.
4.1 Dataset and Preprocessing
OnlyFlow is trained on the WebVid dataset [1], which comprises of 10.7M video-caption pairs, totaling 52K hours of video content. The majority of these videos have a length between a few seconds and 1min, with 100k to 500k pixels by frame. All videos are spatially resized to $256\times 384$ pixels, and $F=16$ consecutive frames are randomly chosen, matching the training practice used in AnimatedDiff [14].
4.2 Model training and inference
Diffusion loss and training pipeline
Following the original formulation of video diffusion model training, the loss function is adjusted to include our control signal, i.e the optical flow $\mathbf{F}=\mathcal{T}(\mathbf{V})\in\mathbb{R}^{F\times H\times W\times 2}$ of the auxiliary input video $\mathbf{V}$. The optimization objective minimizes the mean squared error between the predicted noise and the actual noise:
$\mathcal{L}\left(\theta_{1},\theta_{2}\right)=\mathbb{E}_{\mathbf{z}_{0},\epsilon,c,t}\left[\left|\epsilon-\epsilon_{\theta_{1}}(\mathbf{z}_{t},p,\Phi_{\theta_{2}}(\mathbf{F}),t)\right|_{2}^{2}\right]$ (3)
where $\epsilon_{\theta_{1}}$ is conditioned on $c=\{p,\mathbf{F}\}$ containing the text prompt $p$ and the optical flow $\mathbf{F}$ with the use of a jointly trained encoder $\Phi_{\theta_{2}}.$
During the training phase, we freeze the image generation U-Net and the previously added temporal layers of AnimateDiff. Only the optical flow encoder and the merging linear layers in the attention blocks are trained, using an optical flow conditioning strength $\gamma=1.0$. To mitigate the distribution shift between the WebVid dataset and the dataset used to train the base T2I U-Net, we incorporate the domain adapter used in AnimateDiff training. This adapter is implemented as a LoRA [21] applied to the U-Net spatial layers.
Optical Flow extraction
We employed the commonly used RAFT-Large [33] model to generate a flow field $\mathbf{F}$, with the default values in RAFT inference settings. RAFT is a classic optical flow estimation model, it is fast during inference and widely used for computer vision tasks.
The base T2I model used in this work is the common Stable Diffusion 1.5 model composed of 4 block in its U-Net encoding and decoding part. Consequently, we use an encoder designed to output a corresponding number of optical flow conditioning signals $\Phi\left(\mathbf{F}\right)=\{C_{1},C_{2},C_{3},C_{4}\}$.
We opt to always insert the optical flow conditioning representations into the temporal attention blocks. This decision is motivated by the temporal attention layer’s ability to capture temporal relationships, which aligns with the sequential and causal nature of optical flow.
Spatial and temporal consistency
Nevertheless, it also makes sense to inject the features into the spatial attention layers. The spatial attention layers of the U-Net might adapt the generated visuals to the distortion brought by the optical flow conditioning. Thus, we also study a joint injection of features in both spatial and temporal attention layers, and we label this variation OnlyFlow-ST, while the variation with only temporal attention is labeled OnlyFlow-T.
Learning setup
We train the model for approximately 20 hours using 8 A100 GPUs. The training process involved a batch size of $32$, a constant learning rate of $1\times 10^{-4}$ with the Adam optimizer. The model consists of 198M parameters in the optical flow encoder $\Phi$ and 19M for the injection layers in the U-Net attention in OnlyFlow-T (resp. 31M for OnlyFlow-ST variant).
Using the trained model
During inference, we can use our model by sample noise in the latent with a text prompt and an auxiliary video. It generates a video whose content best represents the content of the text prompt, while following the dynamics of the motion of the auxiliary video. The two input conditions interact via Eq. (2), where the $\gamma$ hyper-parameter manages the influence of each. Our training was carried out for $\gamma=1.0$, so this is the nominal value for using our model in inference. However, it is possible to vary this parameter between 0 (the optical flow is not taken into account) and 1 to play with the progressive inclusion of motion. This could also enable us to investigate how our model works, by offering comparative videos according to $\gamma$ strength.
With our framework, we can process videos of up to 16 frames. However, there are a number of recent proposals [27, 35], to generate an arbitrary long video without any additional training costs.
5 Experimental Results
5.1 Quantitative results
0.00.20.40.60.81.0Optical flow scale in attention3000350040004500FVDOnlyFlow-TOnlyFlow-ST
(a) Fréchet Video Distance (lower is better) between source generated video.
0.00.20.40.60.81.0Optical flow scale in attention1234optical flow absolute mean difference
(b) Mean absolute difference in optical flow between source and generated video.
0.00.20.40.60.81.0Optical flow scale in attention29.429.629.830.030.2CLIP Score
(c) CLIP score similarity between prompts and generated video.
Figure 4: Metrics computed on OnlyFlow-generated videos for both feature maps injection strategies (OnlyFlowT (orange) and OnlyFlowST (blue).) for different values of $\gamma$. The motion realism depicted by fig.(a) improves with $\gamma$, as well as the optical flow similarity in fig.(b). Fig.(c) shows us that the CLIP score does not deteriorate and therefore that the video respect the prompt just as well
Experimentation Goals
As a quick recap, our main goal is to transfer the motion cues from a video $\mathbf{V}$ to the generated one $\mathbf{\hat{V}}$ while remaining realistic and faithful to the textual prompt. Therefore, we assess our model quantitatively using three main components: realism, flow fidelity, and textual alignment.
First, to ensure that our model produces realistic videos, we report the mean Fréchet Video Distance (FVD) [34] across frames. This metric evaluates the temporal coherence of generated videos compared to real ones from a distributional perspective. The FVD between a set of real and generated videos, $R$ and $G$, respectively, is given by
$\begin{split}FV&D(R,G)=\\ &|\mu_{R}-\mu_{G}|+\text{Tr}\left(\Sigma_{R}+\Sigma_{G}-2(\Sigma_{R}\Sigma_{G})^{\frac{1}{2}}\right),\end{split}$ (4)
where $\mu_{I}$ and $\Sigma_{I}$ are the mean and covariance of the video encodings in set $I\in{R,G}$. We follow standard practices proposed by Unterthiner et al. [34] to compute the FVD.
Next, we assess the flow fidelity between the generated video and the prompted one, i.e. whether the produced video contains the same motion patterns as the input. To evaluate this, we measure the absolute pixel-wise difference between the optical flow of the input video and the generated data. The error measurement is given by
$\mathbb{E}_{\mathbf{V},F,H,W}\left[\left|\mathcal{T}(\mathbf{V})-\mathcal{T}(\hat{\mathbf{V}})\right|\right],$ (5)
where $\mathcal{T}$ is the optical flow extractor, RAFT in this case.
Finally, to assess the textual alignment between the input text and the generated video, we adopt the CLIP similarity (CS). Specifically, we focus on the expected CS across frames, as described in the next equation:
$\text{CS}\left(V,p\right)=\mathbb{E}_{F}\left[max\left(100\cos\left(E_{\hat{\mathbf{V}}},E_{t}\right),0\right)\right],$ (6)
where $E_{t}$ is the textual CLIP embedding, while $E_{\hat{\mathbf{V}}}$ is the CLIP vision embedding of the generated video. To study the CS, we used a OpenAI Vision Transformer (ViT) architecture as the image encoder, with a patch size of 16x16 pixels [28].
Empirical Results
To empirically validate OnlyFlow, we perform extensive testing using a subset of 70,000 unseen videos from the WebVid [1] dataset to condition our proposed model. The result are shown in Fig. 4. For a fair comparison with standard T2V generative models, we generate videos with $F=16$ frames at a resolution of $512\times 512$ pixels. In addition, to analyze the impact of our proposed motion module, we evaluate different optical flow conditioning strengths $\gamma$ (Eq. (2)).
First, we discuss the FVD metric in Fig.4(a). Interestingly, we notice that as we increase optical flow conditioning strength $\gamma$, the FVD decreases. We posit that this is explained by the addition of real data as a condition to video generation. Real data brings realistic motion which will end up in being used by OnlyFlow, which in turn will be interpreted as accurate by the FVD metric. In this generative process, enforcing the optical flow of an input video with realistic motion thus naturally improves FVD.
Next, we review the motion fidelity in Fig.4(b), measured as the absolute difference in optical flow. As expected, the difference between optical flow of the input video and optical flow of the generated video decreases as the optical flow conditioning strength increases. This finding implies that the optical flow of the generated video resembles the one of the source video, effectively showing that OnlyFlow achieves our main objective of transferring the motion cues.
Finally, we examine the textual alignment results in Fig. 4(c). The outcome of this experiment show that the model produces stable CLIP scores when using any level of optical flow conditioning. This suggests that the content generated by the model is aligned with the text to the same extent, whether the motion cues are used or not.
5.2 Qualitative results
We now assess OnlyFlow from a qualitative perspective and compare them with previous works. To do so, we present different generative results using the OnlyFlow-T version.
Optical flow conditioning strength
Our model can be utilized with various values of conditioning strengths $\gamma$ to control of the movement in the generated video. In Fig. 5, we present an illustration of this phenomenon using the fixed textual prompt Wind turbine in a field, cloudy weather. In this particular example, no rotative movement is observed when OnlyFlow is disabled ($\gamma=0$). As the conditioning strength increases, the rotative axes align, followed by the blade shape and length, with small wind turbines appearing in the background on tall grass in the auxiliary video. Manually tuning $\gamma$ during the inference can help the user to better express the desired motion.
Figure 5: Illustration of OnlyFlow’s optical flow conditioning strength impact on generated videos. Setting the $\gamma$ scale at which the features from the input video are injected sets the influence of its motion on the generated video. The videos presented in each row are obtained for increasing values of $\gamma$ between 00 and $1.0$. We observe a progressive alignment of the generated video motion and wind turbine position on the auxiliary video
Figure 6: Qualitative comparison of video-to-video generation models. Videos are generated using same text prompt and input video. OnlyFlow exhibits a superior combination of motion fidelity and image realism. It positively compares to approaches that uses depth map (RAVE, VideoComposer, Control-A-Video), is comparable to Gen-1 [12]’s temporal coherence and VideoComposer [37]’s image quality.
Comparison with state-of-the-art video-to-video generation models
For the loosely-defined problem of video-to-video editing and customization, we show that our OnlyFlow model performs better for tasks that require only the motion of the input video. The visual comparison in Fig. 6 reveals that OnlyFlow can generate animations whose movements are close to the input video, given a different textual prompt. In this comparison, the rippling waves effect of the auxiliary video struggle to be accurately represented in other models, except for Control-A-Video and Gen-1. We notice that the prompt adherence is also better followed in our approach, Gen-1, VideoComposer and Tune-A-Video. Nevertheless, the latter two models did not achieve to take motion inspiration from our auxiliary video and Gen-1 seems to stack mountains and ripples visuals. Compared to other work such as Tune-A-Video [39], our approach is lightweight. On top of that, it does not require any costly DDIM [32] inversion process used in TokenFlow [13] and RAVE [23].
Comparison with modern camera-movement controlled video generation
Figure 7: Qualitative comparison on camera-movement controlled video generation. All videos are generated using the same text prompt, camera trajectory (pan-left), and seed when available. Without having trained for this task, our OnlyFlow achieve the same camera control capability as other camera-movement approaches trained on this task.
In recent work, such as MotionCtrl [38] and CameraCtrl [17], the conditioning of camera motion is achieved with a camera position encoding, while the AnimateDiff [14] motion conditioning uses a LoRA trained on the temporal layer for a given motion. For camera displacements, OnlyFlow provides a flexible alternative for modeling camera movement. To do this, we create an optical flow to describe the desired motion or provide a video with the desired motion. In Fig. 7 we show a specific scenario where we prompted our model with a constant horizontal movement to the left. The generated results clearly show that the generation is on par with the controllability of specific approaches such as CameraCtrl or MotionCtrl. Note that we did not observe satisfactory results with AnimateDiff’s motion LoRA, both in terms of aesthetics and expected camera movements. This shows the versatility of our model.
Semantic alignment
We noticed an interesting capability of semantic alignment between the descriptive information contained in the optical flow fed into the model and the given prompt. As shown in Fig. 8, the optical flow inadvertently captures object shapes. When there is a match between the input video motion concept and the prompt, the model successfully maps the zones of interest to be generated. This ability is similar to other conditioning models such as ControlNet [43] and T2I-Adapter [25], but our method does not require manual control map creation.
(a) Frames from the auxiliary video of which optical flow is extracted
(b) Frames from corresponding generated video
Figure 8: Example of semantic alignment between optical flow and prompt. With the prompt ”Trees in forest”, our OnlyFlow model generates a clothesline corresponding to the teeth of the smile.
User preference study
We employ the Two-alternative Forced Choice (2AFC) protocol for text-driven video editing [2, 11, 13, 26, 40]. Participants are presented with the input video, our results and a baseline, and are asked to determine which video better aligns with the text prompt, which video better preserves the motion of the input video, and which video they prefer overall. We collected 300 user judgments. As seen in Tab.1, our method is consistently preferred over all baselines.
Ref. Method Judgements in our favor (%)
TokenFlow [13] 54.5%
Control-A-Video 71.1%
RAVE [23] 62.1%
Gen-1 [12] 60.6%
VideoComposer [37] 63.6%
Table 1: User study: preference (%) between OnlyFlow and reference methods. We report the percentage of judgments in favor of OnlyFloww.r.t. each baseline.
color=yellow!60,inline]H:
Parler des results des 2 types de train
Quantitatif: - diff optical flow - FVD - Clip Score
- CLIP similarity among consecutive video frames (optionnel)
- courbe de training (optionnel) ou annexe
Figure belle pour montrer le comportement de OnlyFlow et que ça marche bien
- flow strength figure image ¿¿ exemple random de comment OnlyFlow marche (avec un exemple juste joli, rien de spécial aka exemple canonique (a remettre dans la figure d’intro du papier)) (2/3 frames + 5 valeurs strength + 2/3 vidéos)
Figure fonctionnalités out of the box
Comparaisons qualitatites avec autres algos:
- CameraCtrl ex avoir des pan ou certains type de mouvements (on compare cet algo avec le fait de conditionner OnlyFlow par des videos de pan et on montre que OnlyFlow sait faire du cameractrl, sans pouvoir avoir été entrainé que à faire ça (point positif))
(comp algo MotionCtrl, CameraCtrl, AD LORA)
- ControlNet ex avoir idem mais ce qu’il peut faire ou non (on compare cet algo avec le fait de conditionner OnlyFlow par des vidéos de type mask que l’on donnerait à controlnet, et on montre que OnlyFlow sait faire du ctrlnet, sans pour autant avoir été entrainé à en faire).
Figure pour montrer que OnlyFlow s’interface bien avec les autres types de conditionnement
[[[[ sparse ctrl, ipadapter, lora ]]]]
Montrer: - ok style transfer - ok pour tache type ControlNet - contrôle de camera
color=teal!60,inline]M: Messages principaux figures resultats: - figure clipscore : ca part pas en vrille: ca ne deteriore pas le suivi du prompt - diff optical flow cut à 0 et baisse - FVD : ca baisse un peu: donc plus fidele, amelioration
6 Discussion and limitations
color=teal!60,inline]M: - lightweight - plus pour gratuits par rapport à - generalisation de controlnet et camera control dans un framework unique (meme si tâche specifique, ex : cameractrl, MotionCtrl)
Discussion
Another motivation for OnlyFlow is the creation of artistic videos. Using the natural motion present in video footage allows the generation of videos that would require tremendous editing efforts to be created without OnlyFlow. In the supplementary video, we provide examples of artistic videos we created during the project, that incorporate upscaling with ESRGAN [36] and frame interpolation with FILM [29], which by themselves showcase the type of outputs that can be done using OnlyFlow. Moreover, OnlyFlow is readily compatible with extensions of AnimateDiff and Stable-Diffusion that are often used in artistic creation context such as IP-Adapter [41], which can further be used to improve the quality and controllability of the generated videos.
Limitations
One clear limitation of our work is photorealism obtained in the generated videos. While T2V models like AnimateDiff [14] are powerful generative models, their generation remain poor for most complex prompts depicting scenes with various objects and subjects or scenes with movements that might not appear in the training set. The resolution of the generated video are also limited by the underlying model.
Another limitation is that while optical flow as motion conditioning is useful and efficient in most cases as shown in our experiments, optical flow is not the only way to estimate motion in a video. We could use other motion conditions. For instance, optical flow does not always allow separate camera movements from object movements. Thus, in order to better model the motion of an input video, other methods could be investigated, such as Kalman Filtering, Block or Feature Matching, Homography-based Motion estimation. A combination of such motion estimations could lead to better results, and are left for future work.
7 Conclusion
In this paper, we introduced our OnlyFlow method, which improves on T2V models such as AnimateDiff by conditioning it on the motion extracted from an input auxiliary video. The optical flow of the auxiliary video is then passed to a trainable flow encoder connected to the T2V model. The flow encoder outputs features maps which are injected into the attention layers of the T2V’s U-Net. The generated video is encouraged to accurately follow the motion of the input video. While other methods allow conditioning by movement (with motion masks or user-defined vectorial camera movements), OnlyFlow allows to generate any video with a text prompt and the motion of an auxiliary video. In our experiments, we demonstrate the efficiency of OnlyFlow for various applications such as camera motion control (panning, zooming, etc.) and video editing/style transfer. Without necessarily being trained for such applications, OnlyFlow compares positively to other state-of-the-art methods that have been specifically trained for such applications. We also carried out a user study which confirmed the interest of our method in generating very interesting videos, demonstrating that their movement convincingly and coherently follows the input reference.